the problem is to complete the table of oredered pairs for the equation and plot the ordered pairs for 3x+2y=6. They gave us for the first one x is 0 secord y is 0, the third x is 1, the fourth was y is 3.
What I came up with was
(o,3) (2,o) (4,3) I am not sure if I did it right.

Answer by checkley77(12844)

x is 0 secord y is 0, the third x is 1, the fourth was y is 3.
3x+2y=6
3*0+2Y=6
0+2Y=6
2Y=6
Y=6/2
Y=3 ANS. (0,3)
3X+2*0=6
3X+0=6
X=6/3
X=2 ANS. (2,0)
3*1+2Y=6
3+2Y=6
2Y=6-3
2Y=3
Y=3/2 ANS. (3/2,2)
3X+2*3=6
3X+6=6
3X=6-6
3X=0
X=0 ANS. (0,3)